Grade 12th passAlgebra

Write the following rational numbers in ascending order:
-1/3, -2/9, -4/3

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

To convert the given rational numbers into like fraction we have to find LCM, LCM of 3, 9, and 3 is 9
Now,
(-1/3)= [(-1 × 3)/ (3 × 9)] = (-3/9)
(-2/9)= [(-2 × 1)/ (9 × 1)] = (-2/9)
(-4/3)= [(-4 × 3)/ (3 × 3)] = (-12/9)
Clearly,
(-12/9) < (-3/9) < (-2/9)
Hence,
(-4/3) < (-1/3) < (-2/9)

Thanks